OpenStudy (waheguru):

Recursive Formula Question t(n) = t(n-1)+12 t(1) = 5 If I were to find term 1 using this formula I should be getting 5 but I think I am getting 13 instead?

OpenStudy (waheguru):

Because t(1) - 1 = t(0) + 12 = 12 ?

OpenStudy (anonymous):

What's t(0) ? Is it negative?

OpenStudy (waheguru):

but then how can we find term 1 using the formula?

OpenStudy (anonymous):

If you don't have t(0) since it's recursive you can use t(2) instead if you have it, and plug 2 in for n, which would let you do it properly. If you only have t(1) then you can't do anything.

OpenStudy (anonymous):

Well obviously you can do that, but if the question is what he worded, he has to prove t(1) works within that equation and he would have to have another term such as t(0) or t(2) to do that proof.

OpenStudy (waheguru):

So @G-unit to find t1 we would need another poitn right? or else its not possible?

OpenStudy (waheguru):

what is discrece math|

OpenStudy (waheguru):

btw I made this question up to have a better understaiding of sequencing so I can understand and do much more complex problems and geometric sequences...

OpenStudy (loser66):

@G-unit please